Actual sales of TV's for the first six month in 2021 were as follows:
January      60            February     80          March   42
April           68            May             44           June     72

Use this information and determine in whole numbers:

1 The average actual monthly sales for that period         

2 Trends: Reduce this average by 6% per month for the next three months   

3 Simple Moving Average (SMA): make a forecast for July using a three months simple moving
average.       


4 Weighted Moving Average (WMA): make forecast for July, using a three months weighted
moving average where the weights are 0.5; 0.3 and 0.2 respectively.        
           

5. Make a forecast for July, using the exponential smoothing technique, if the forecast for June
was 74 and the smoothing constant is 0, 5.                                                                   
Answer #1

1. average actual monthly sales: year 2021

Jan 60

Feb 80

Mar 42

Apr 68

May 44

June 72

Average= (60+80+42+68+44+72)/6= 61

2. Reducing average by 6% for next 3 months  

July= 61-(61*.06)= 57.34

August= 57.34- (57.34*.06)= 53.89

September= 53.89- (53.89*.06)= 50.65

3. Forecast for july using simple moving average:

Demand for last 3 months Months

June     72

May 44

April 68

using 3 months simple moving average:

Demand forecast for 7th month= ( demand for 6th month+ demand for 5th month+ demand for 4th month)/ 3

= (72+44+68)/3

= 61.33

4. 3 months weighted moving average

Demand for last 3 months weights

June 72 0.5

May 44 0.3

April 68 0.2

Weighted average = (72*0.5) + (44*0.3) + (68*0.2)/ (0.5+0.3+0.2)

= (36+13.2+13.6)/ 1= 62.8